Two bicycles tyres are set rolling with the same initial speed of 3.5m/s on a long straight road and the distance each travels before its speed is reduced by half is measured. One tyre is inflated to a pressure of 40 psi and goes 18.1m; the other is at 105 psi and goes 92.9m, what is the coefficient of rolling friction for each? Assume that the net horizontal force is due to rolling friction only.
1
Expert's answer
2020-07-08T13:11:56-0400
"(0.5v)^2-v^2=-2\mu_1 gs"

"0.75(3.5)^2=2(9.8)(18.1)\mu _1\\\mu _1=0.026"

"(0.5v)^2-v^2=-2\mu_2 gd"

"0.75(3.5)^2=2(9.8)(92.9)\mu _2\\\mu _2=0.005"